Circular No. 6250

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1995 S3 (MRKOS)

Reports of CCD recoveries of comet P/1991 F1 (= 1991 IV = 1991k) have been received from C. W. Hergenrother and W. Offutt, as shown below. The indicated correction (ignoring perturbations) to the prediction by S. Nakano on MPC 22033 (ephemeris on MPC 25447) is ΔT = +0.08 day.

1995 UT α_{2000} δ_{2000} m_{1} Observer
Sept.20.42429 3^{h}14^{m}55.^{s}23 +45^{o}29′58′′.7  Larson
20.42696 3 14 55.26 +45 29 59.3  _{”}
20.43059 3 14 54.97 +45 30 02.0  _{”}
20.43899 3 14 54.96 +45 30 07.5  _{”}
28.37233 3 11 15.29 +46 42 59.7  Offutt
28.39525 3 11 14.59 +46 43 05.5  _{”}
Oct. 21.22454 2 49 34.18 +49 29 10.3 20.2  _{”}
21.22968 2 49 33.82 +49 29 10.2 20.3  _{”}
21.23677 2 49 33.19 +49 29 12.3 20.5  _{”}
21.24789 2 49 32.40 +49 29 16.3 20.2  _{”}
21.26616 2 49 31.06 +49 29 21.2 20.5  _{”}
21.29030 2 49 28.83 +49 29 28.8 20.6  _{”}
21.30182 2 49 27.88 +49 29 35.8 19.1  _{”}
21.30848 2 49 27.76 +49 29 35.2 19.2  _{”}

S. M. Larson (Lunar and Planetary Laboratory). 2.3-m f/2 Steward reflector, Kitt Peak. Comet identified and measured by C. W. Hergenrother. Stellar appearance. Red magnitude m_{2} = 21.7 and 22.3 on first two images.

W. Offutt (Cloudcroft, New Mexico). 0.6-m f/7 Ritchey-Chrétien. Signalto-noise very poor on Sept. 28, image fainter than mag 21. Low signalto-noise with broadband skyglow filter on Oct. 21, measurement compromised by proximity to star on Oct. 21.266, 21.302 and 21.308.

The following improved orbital elements, by B. G. Marsden, satisfy 62 observations 1991–1995 with mean residual 1′′.01:

Epoch = 1996 Nov. 13.0 TT

T = 1996 Nov. 9.0771 TT  ω = 180.^{o}5221
e = 0.554249  Ω = 1.6528 }2000.0
q = 1.412898 AU  i = 31.4708

a = 3.169705 AU n^{o} = 0.1746529 P = 5.643 years

1995 October 23 Brian G. Marsden